OpenStudy (kitkat1):

The sides of a square are three to the power of two sevenths inches long. What is the area of the square?

OpenStudy (kitkat1):

nine to the power of four sevenths square inches three to the power of four sevenths square inches three to the power of the fraction four over forty nine square inches nine to the power of the fraction four over forty nine square inches

OpenStudy (campbell_st):

so the side length is \[3^{\frac{2}{7}}\] then the area is \[A = 3^{\frac{2}{7}} \times 3^{\frac{2}{7}}\] so use the indea lat form multiplying the same base.... add the powers \[x^a \times x^b = x^{a + b}\]

OpenStudy (kitkat1):

Im not good at this can u show me how?

OpenStudy (campbell_st):

add you need to do is add the fractions... both numbers have 3 as a base

OpenStudy (kitkat1):

Okay so 3 and 4/14

OpenStudy (kitkat1):

But there is no answer like that

OpenStudy (kitkat1):

So how do I find the answer?

OpenStudy (campbell_st):

no, when you add fractions with the same denominator... just add the tops 5/11 + 2/11 = 7/11 notice the denominator doesn't change

OpenStudy (kitkat1):

So what will be the answer?

OpenStudy (campbell_st):

well go back and add the 2 fractions...

OpenStudy (kitkat1):

thats what I did and it ended up 3 and 4/14

OpenStudy (kitkat1):

Could u show me how to do it Im really bad at this

OpenStudy (campbell_st):

here is a link to a fractions calculator... try it http://www.calculatorsoup.com/calculators/math/adding-fractions-calculator.php

OpenStudy (campbell_st):

or on your calculator 2/7 + 2/7 and see what you get

OpenStudy (kitkat1):

And?

OpenStudy (kitkat1):

Im stuck I dont know what to do

OpenStudy (campbell_st):

just add the numerators... and leave the denominator as 7

OpenStudy (kitkat1):

4?

OpenStudy (campbell_st):

so using your idea if you have 1/2 cheese pizza another 1/2 of mushroom pizza you only have 1/4 of a pizza

OpenStudy (campbell_st):

so 4 over what...?

OpenStudy (kitkat1):

7

OpenStudy (kitkat1):

3 and 4 over 7

OpenStudy (campbell_st):

there you go... that's the answer